# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v2

@@ Abstract
The rapid advances of multimodal agents built on large foundation models have largely overlooked their potential for language-based communication between agents in collaborative tasks. This oversight presents a critical gap in understanding their effectiveness in real-world deployments, particularly when communicating with humans. Existing agentic benchmarks fail to address key aspects of inter-agent communication and collaboration, particularly in scenarios where agents have unequal access to information and must work together to achieve tasks beyond the scope of individual capabilities. To fill this gap, we introduce a novel benchmark designed to evaluate the collaborative performance of multimodal multi-agent systems through language communication. Our benchmark features a variety of scenarios, providing a comprehensive evaluation across four key categories of agentic capability in a communicative collaboration setting. By testing both agent-agent and agent-human collaborations using open-source and closed-source models, our findings reveal surprising weaknesses in state-of-the-art models, including proprietary models like GPT-4o. Some of these models struggle to outperform even a simple random agent baseline in agent-agent collaboration and only surpass the random baseline when a human is involved.
Machine Learning, ICML
@@ 1 Introduction
The field of multimodal agents is experiencing rapid growth  (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024), with research efforts expanding at an unprecedented pace. However, amidst this growth, a critical gap in research has emerged: the lack of focus on collaborative work (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b) among multiple multimodal agents. Synergistic operation of such agents is a highly promising but largely unexplored domain. Language agents can collaboratively finish complex tasks such as software development  (Qian et al., 2024; Du et al., 2024) or even machine learning research  (Schmidgall et al., 2025) by assuming functional roles such as system designer, function generator, etc. Current research on multimodal agents (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024) has mainly focused on individual agent capabilities, neglecting the potential for inter-agent collaboration. This limitation is further compounded by existing benchmarks such as TheAgentCompany (Xu et al., 2024a), VisualWebArena (Koh et al., 2024) and MME-RealWorld (Zhang et al., 2024), which do not assess collaborative performance between agents. As a result, our ability to evaluate and improve multi-agent systems remains constrained, hindering progress in this crucial area.
@@
Figure 1: Overview of the interaction between the Solver and Expert agents in our benchmark. Both agents operate with structured input corresponding to working and episodic memory. The Solver receives an image of the puzzle state (working memory) and makes decisions based on the available actions described in the task prompt. The Expert, guided by instruction manuals (working memory), provides advice based on the Solver’s descriptions, such as indicating which buttons to press. The Solver can choose to execute actions by interacting with the environment or communicate with the Expert for further guidance. Their interaction is documented through a dialogue, showcasing the cooperation required to complete the task. Both agents engage in self-reflection by referencing the conversation history, which is continuously updated and incorporated into their input as episodic memory.
Motivated by these aforementioned issues, we propose a novel benchmark for evaluating collaborative multimodal multi-agent frameworks to address critical gaps in current approaches (see  Figure 1). Our evaluation setting also simulates a scenario where an in-house agent with direct access to sensitive data (i.e., the AI solver) collaborates with external expert agents (i.e., the AI expert) to analyze information without compromising privacy. This evaluation setting simulates how we handle and extract insights from sensitive datasets across various domains.
We assess multimodal multi-agent systems using a series of carefully designed collaborative puzzle games. Building on cognitive science research employing simple puzzles to assess cognitive ability  (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006), we design vision-language puzzles to evaluate the core cognitive abilities of multimodal agents, including learning, adaptation, and problem-solving. These scenarios typically involve two-player setups where agents have access to different, complementary information. (i.e., in a bomb defusal game, one agent possesses details about the bomb, while the other has access to a disarming manual). By employing such diverse and interactive scenarios, we aim to provide a thorough assessment of multimodal multi-agent performance.
Our benchmark includes 10 distinct, easily customizable puzzles with thousands of unique solutions. We tested two different settings (AI-AI and AI-Human) and evaluated several popular multimodal models, including closed-source models (GPT-4V  (Achiam et al., 2023), GPT-4o  (Hurst et al., 2024)) and open-source models (QwenVL (Bai et al., 2023), and InternVL  (Chen et al., 2024)). Surprisingly, even powerful closed-source models such as the GPT series do not outperform a human solver, highlighting a potential growth area for future model development. Our contributions are as follows:
  • We propose an evaluation framework called COMMA, a multimodal agent benchmark focusing on language communication between multiple agents (Section 3).
  • Using COMMA, we carefully record conversations and performance metrics between state-of-the-art multimodal models such as QwenVL, InternVL, LLaMA, Gemini, GPT-4o, etc (Section 4).
  • We categorize the agent capabilities tested in our model and common failure modes, providing insight into potential future research directions for improving inter-agent communication (Section 5).
@@ 2 Related Work
Instruction-based agent benchmarks evaluate an agent’s capability of following a human instructions to finish a task (e.g., navigating on a website, interacting with an operating system  (Xu et al., 2024b; Xie et al., 2024; Cao et al., 2024)). However, our benchmark focuses more on a communication-based evaluation where two clients engage in multi-turn conversations to solve a task collaboratively.
@@ 3.1 Design Principles of the Benchmark
Our benchmark is inspired by the cooperative gameplay scenario in Keep Talking and Nobody Explodes (Games, 2015). In this game, two players work together to defuse a bomb under time pressure. One player, the defuser, can see the bomb but lacks the instructions to disarm it. The other player, the expert, has access to the bomb’s manual but cannot see the bomb itself. The players must rely on effective communication to exchange information, navigate challenges, and defuse the bomb.
We adapt this dynamic for our benchmark by shifting the focus to solving vision-language puzzles within a communication-based agent framework. To better reflect this broader scope, we rename the defuser role to “Solver,” emphasizing its general-purpose functionality across diverse tasks. As multimodal agent systems continue to gain traction, our benchmark aims to rigorously evaluate their reasoning, communication, and collaborative capabilities. Its design is grounded in the following core principles:
@@
We carefully structure the Solver and Expert agents, drawing on the cognitive agent terminology from  Sumers et al. (2023). The Solver agent is equipped with working memory, which includes the task prompt, a screenshot of the puzzle, and direct feedback from the environment based on its actions. Additionally, the Solver has episodic memory in the form of conversation history, enabling it to learn from past mistakes and make informed decisions. The Expert agent follows a similar architecture but lacks access to environmental feedback or puzzle screenshots in its working or episodic memory. Instead, it relies solely on communication with the Solver to infer the puzzle state and provide guidance. Figure 1 provides an illustration of the interaction between the agents.
Intelligence Testing in Cognitive Science:
Our benchmark draws inspiration from the foundational principles of intelligence, often defined as the ability to learn from experience, adapt to the environment, and solve problems using cognitive skills (Kempf-Leonard, 2005). Cognitive science research has shown that even simple tests can effectively measure cognitive ability (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006). Standardized intelligence tests, such as MENSA (MENSA International, n.d.) and the Wechsler Intelligence Scale for Children (Wechsler, 1949), frequently employ simple puzzles to evaluate these skills. Building on this approach, our benchmark aims to assess the core cognitive capabilities of multimodal agents by creating simple vision-language puzzles tailored to test these abilities.
@@
3.2 Categories of Agent Capability
We benchmark agents working under different roles to solve various tasks in multiple settings, each requiring different capabilities. Specifically, the Solver agent must demonstrate strong instruction-following and multimodal reasoning, while the Expert agent is expected to excel in long text summarization and information retrieval. Both agents must possess visual comprehension and descriptive skills to succeed. Below, we outline the core capabilities tested in our benchmark.
@@ 3.3 Tasks
  • ColorPuzzle ( MR , MSR ): The solver aims to turn all of the squares in a 4x4 grid white. At each step, the solver should press squares based on the frequencies of colors, following the rules specified in a table.
@@ 4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. We use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle and run inference on a single NVIDIA A100 GPU with 80GB RAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
@@ 4.2 Evaluation Metrics
4.3 Models
Open-Source Models
  • QwenVL  (Bai et al., 2023): We use QwenVL-2, offering enhanced pretraining for improved performance on vision-language tasks. We use the 7b variant.
  • LLaVA  (Liu et al., 2024a): We use version 1.6 of LLaVA with a Mistral 7b language model backbone.
Closed-Source Models
  • GPT-4V  (Achiam et al., 2023): A version of OpenAI’s GPT-4, GPT-4V incorporates visual processing, enabling it to interpret both text and images.
  • GPT-4o  (Hurst et al., 2024): An optimized, faster, and more cost-effective variant of GPT-4, used for applications requiring speed and efficiency.
  • Gemini 2.0  (Anil et al., 2023): Google’s most recent LLM with a focus on agent capabilities supporting different input modalities such as vision, audio and text.
@@ 5.1 Overall Performance
Table 1 presents the average partial success rate (%) of various multimodal agents and a human solver across several puzzles, highlighting their relative performance. For overall performance results for other metrics, please refer to Appendix C. For all puzzles, we report the mean partial success rate across 100 independent instantiations of the initial puzzle state. Due to the source of randomness being the puzzle initialization, rather than differences in model output on the same data, we report standard error estimates for all puzzles (±$\pm\sigma$) to approximate the confidence interval of performance on each puzzle.
@@
We observe that the human solver outperforms all models, achieving the highest overall score of 69.01%. Among the AI models, GPT-4o demonstrates the best overall performance (41.74%), significantly surpassing the others, including Gemini-2.0 (33.37%) and GPT-4V (27.62%). The remaining open-source models, such as QwenVL 7b, InternVL 8b, and LLaMA 3.2, show significantly lower success rates, with LLaVA-1.6 achieving the lowest overall score (17.32%). Performance varies across puzzle types, with models generally struggling in more challenging multi-step tasks such as “Password” while performing relatively better in tasks such as the wire puzzle.
5.2 Qualitative Analysis on Model Failures
Model Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 100 ± 0.0 65 ± 15.0 90 ± 10.0 30 ± 13.3 50 ± 16.7 60 ± 11.9 80 ± 13.3 17 ± 5.7 70 ± 15.1 100 ± 0.0 69.01 ± 4.2 GPT-4o 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 Gemini 85 ± 3.6 47 ± 4.9 35 ± 4.8 60 ± 3.3 24 ± 2.0 24 3.3 4 ± 2.0 11 ± 1.3 16 ± 3.3 27 ± 4.5 33.37 ± 1.3 GPT-4V 77 ± 4.2 48 ± 5.0 39 ± 4.9 30 ± 3.0 8 ± 1.4 32 ± 2.9 0 ± 0.0 8 ± 1.1 15 ± 3.5 19 ± 3.9 27.62 ± 1.3 LLaMA 3.2 64 ± 4.8 13 ± 2.5 27 ± 4.5 29 ± 3.5 27 ± 3.0 28 ± 2.9 0 ± 0.0 9 ± 1.4 24 ± 3.2 0 ± 0.0 22.15 ± 1.1 QwenVL 56 ± 5.0 40 ± 4.9 26 ± 4.4 31 ± 3.5 16 ± 2.3 25 ± 2.5 0 ± 0.0 5 ± 0.8 4 ± 1.3 0 ± 0.0 20.28 ± 1.1 InternVL 61 ± 4.9 23 ± 2.5 28 ± 4.5 25 ± 3.3 18 ± 1.4 24 ± 2.5 1 ± 1.0 9 ± 1.4 24 ± 4.7 0 ± 0.0 19.57 ± 1.0 Random 57 ± 5.0 16 ± 2.6 44 ± 5.0 32 ± 3.6 15 ± 1.6 18 ± 2.1 0 ± 0.0 6 ± 1.0 0 ± 0.0 0 ± 0.0 18.70 ± 1.1 LLaVA 1.6 41 ± 4.9 20 ± 2.6 25 ± 4.3 35 ± 3.6 13 ± 1.5 16 ± 2.4 0 ± 0.0 2 ± 0.8 20 ± 4.2 0 ± 0.0 17.32 ± 1.0
Table 1: Average partial success rate of multimodal agents on each puzzle with standard error estimates for uncertainty ±$\pm\sigma$. For each row, the solver and expert are separate instances of the same model. “Human” model indicates a human is the solver, and the expert is a GPT-4o model. The solver is assigned a value between 0-100 indicating how far the solver progressed through the puzzle. The partial success rate is calculated by averaging this value over 100 or 10 independent runs of each puzzle for AI or human solver respectively. The overall column is calculated by averaging across all the puzzles. Bolded values indicate the best partial success rate for the column.
@@ 5.2 Qualitative Analysis on Model Failures
Figure 2: Case study examples of InternVL 8b (bottom left) and GPT-4o (all others) failures when used as agents in our benchmark. Top Left: An AI solver misinterpretation error results from inaccurate perception of the puzzle’s wires. Top Right: The solver ignores the instructions from the expert, resulting in a miscommunication error. Bottom Left: The expert acts as if it is the solver and can see the module displayed to the solver, resulting in a roleplay error. Bottom Right: The solver performs the same action despite being in the same situation seen previously, leading to a repetition loop error.
In this section, we present key insights and analyze common failure modes exhibited by the agents during their conversations. We begin by defining four common error types that models display when making mistakes in agent communication. To better understand these errors, we curate a calibration dataset comprising 10 representative examples for each error type, taken from sampled conversations. This dataset is used to validate and develop a GPT-4o1  (Jaech et al., 2024) model judge. After making minor adjustments to the judge’s prompt and providing access to the ground truth conversations, the model achieved 94% accuracy on the calibration dataset, demonstrating strong alignment with human annotators. Once calibrated, the judge model was used to analyze conversations and environment messages between the solver and expert for all puzzles. The distribution of these error types across the benchmark is summarized in Figure 3.
@@
Figure 3: Distribution of error categories across conversations evaluated by our calibrated GPT-Judge for GPT-4o conversations (left) and LLaMA 3.2 conversations (right).

### v3

@@ Abstract
The rapid advances of multimodal agents built on large foundation models have largely overlooked their potential for language-based communication between agents in collaborative tasks. This oversight presents a critical gap in understanding their effectiveness in real-world deployments, particularly when communicating with humans. Existing agentic benchmarks fail to address key aspects of inter-agent communication and collaboration, particularly in scenarios where agents have unequal access to information and must work together to achieve tasks beyond the scope of individual capabilities. To fill this gap, we introduce COMMA: a novel puzzle benchmark designed to evaluate the collaborative performance of multimodal multi-agent systems through language communication. Our benchmark features a variety of multimodal puzzles, providing a comprehensive evaluation across four key categories of agentic capability in a communicative collaboration setting. Our findings reveal surprising weaknesses in state-of-the-art models, including strong proprietary models like GPT-4o and reasoning models like o4-mini. Many chain of thought reasoning models such as R1-Onevision and LLaVA-CoT struggle to outperform even a random baseline in agent-agent collaboration, indicating a potential growth area in their communication abilities.  [^1] Our data and code is available at https://github.com/tossowski/COMMA
@@ 1 Introduction
The field of multimodal agents is experiencing rapid growth  (Xu et al., 2024c; Xie et al., 2024; Cao et al., 2024), with research efforts expanding at an unprecedented pace. However, amidst this growth, a critical gap in research has emerged: the lack of focus on collaborative work (Gurcan, 2024; Park et al., 2023; Hong et al., 2024; Liu et al., 2024b) among multiple multimodal agents. Synergistic operation of such agents is a highly promising but largely unexplored domain. Language agents can collaboratively finish complex tasks such as software development  (Qian et al., 2024; Du et al., 2024) or even machine learning research  (Schmidgall et al., 2025) by assuming functional roles such as system designer, function generator, etc. Current research on multimodal agents (Xu et al., 2024c; Xie et al., 2024; Cao et al., 2024) has mainly focused on individual agent capabilities, neglecting the potential for inter-agent collaboration. This limitation is further compounded by existing benchmarks such as TheAgentCompany (Xu et al., 2024a), VisualWebArena (Koh et al., 2024) and MME-RealWorld (Zhang et al., 2024), which do not assess collaborative performance between agents. As a result, our ability to evaluate and improve multi-agent systems remains constrained, hindering progress in this crucial area.
@@
Figure 1: Overview of the interaction between the Solver and Expert agents in our benchmark. Both agents operate with structured input corresponding to working and episodic memory. The Solver receives an image of the puzzle state (working memory) and makes decisions based on the available actions described in the task prompt. The Expert, guided by instruction manuals (working memory), provides advice based on the Solver’s descriptions, such as indicating which buttons to press. The Solver can choose to execute actions by interacting with the environment or communicate with the Expert for further guidance. Their interaction is documented through a dialogue, showcasing the cooperation required to complete the task. Both agents engage in self-reflection by referencing the conversation history, which is continuously updated and included in their input as episodic memory. We include a more comprehensive example in Appendix A
Motivated by these aforementioned issues, we propose a benchmark for evaluating multimodal models in a collaborative agentic setting to address critical gaps in current approaches (see  Figure 1). While recent work has explored scaling test time compute  (Snell et al., 2025) to improve reasoning performance, such methods often rely on the unrealistic assumption that a model has sufficient capability and access to all necessary information to complete the task. In contrast, our benchmark poses a unique challenge: models must dynamically alternate between intensive reasoning based on the information available and effective communication to request additional details when needed. Our evaluation setting also simulates a scenario where an in-house agent with direct access to sensitive data (i.e., the AI solver) collaborates with external expert agents (i.e., the AI expert) to analyze information without compromising privacy. Specifically, this work makes the following contributions:
  • We propose an evaluation benchmark called COMMA, a multimodal agent benchmark which uses puzzles to test language communication between multiple agents (Section 3). To achieve a high score, agents must engage in concise, but meaningful dialogue to arrive at each solution.
  • Using COMMA, we carefully record conversations and performance metrics between state-of-the-art multimodal models such as o4-mini, GPT-4o, Gemini, QwenVL, etc (Section 4).
  • We categorize the agent capabilities tested in our model and common failure modes, providing insight into future research directions for improving inter-agent communication (Section 5).
@@ 2 Related Work
Instruction-based agent benchmarks evaluate an agent’s capability of following a human instructions to finish a task (e.g., navigating on a website, interacting with an operating system  (Xu et al., 2024c; Xie et al., 2024; Cao et al., 2024)). However, our benchmark focuses more on a communication-based evaluation where two clients engage in multi-turn conversations to solve a task collaboratively.
Cognitive Science
Our benchmark draws inspiration from the foundational principles of intelligence, often defined as the ability to learn from experience, adapt to the environment, and solve problems using cognitive skills (Kempf-Leonard, 2005). Existing cognitive science research has shown that even simple tests can effectively measure cognitive ability (Davidson et al., 2006; St Clair-Thompson & Gathercole, 2006). Standardized intelligence tests, such as MENSA (MENSA International, n.d.) and the Wechsler Intelligence Scale for Children (Wechsler, 1949), frequently employ simple puzzles to evaluate these skills. Our work extends this research to multi-agent communication, designing vision-language puzzles to evaluate the core cognitive abilities of multimodal agents.
@@ 3.1 Design Principles of the Benchmark
Our benchmark is inspired by the cooperative gameplay scenario in Keep Talking and Nobody Explodes Games (2015). In this game, two players work together to defuse a bomb under time pressure. One player, the defuser, can see the bomb but lacks the instructions to disarm it. The other player, the expert, has access to the bomb’s manual but cannot see the bomb itself. The players must rely on effective communication to exchange information, navigate challenges, and defuse the bomb.
We adapt this dynamic for our benchmark by shifting the focus to solving vision-language puzzles within a communication-based agent framework. While some manuals are similar, we change most of the manuals from the original game to provide a fair comparison between models and prevent data leakage. To reflect this broader scope, we rename the defuser role to “Solver,” emphasizing its general-purpose functionality across diverse tasks. In particular, our benchmark simulates real-world tasks like pair programming with coding copilots and multimodal tutoring systems, where one agent may have access to a visual context and another only to language instructions. Additionally, this setting closely resembles multi-agent workflows involving external tool use, such as querying databases, invoking APIs, or browsing the web. However, we abstract away tool interfaces to focus on evaluation of reasoning, communication, and collaborative capabilities. Our benchmark design is grounded in the following core principles:
@@
We carefully structure the Solver and Expert agents, drawing on the cognitive agent terminology from  Sumers et al. (2023). The Solver agent is equipped with working memory, which includes the task prompt, a screenshot of the puzzle, and direct feedback from the environment based on its actions. Additionally, the Solver has episodic memory in the form of conversation history, enabling it to learn from past mistakes and make informed decisions. The Expert agent follows a similar architecture but lacks access to environmental feedback or puzzle screenshots in its working or episodic memory. Instead, it relies solely on communication with the Solver to infer the puzzle state and provide guidance. We also offer the option for a human to roleplay as either the solver or expert agent in our framework. Figure 1 provides an illustration of the interaction between these agents, and we provide more details in Appendix A.
@@
3.2 Categories of Agent Cognitive Capability
We benchmark agents working under different roles to solve various tasks in multiple settings, each requiring different cognitive capabilities. Specifically, the Solver agent must demonstrate strong instruction-following and multimodal reasoning, while the Expert agent is expected to excel in long text summarization and information retrieval. Both agents must possess visual comprehension and descriptive skills to succeed. Below, we outline the core capabilities tested in our benchmark.
@@ 3.3 Tasks
  • ColorPuzzle ( MR , MSR ): The solver aims to turn all of the squares in a 4x4 grid white. At each step, the solver should press squares based on the frequencies of colors, following the rules specified in a table visible only to the expert.
@@ 4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. For most models, we use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle. However, for reasoning models we set the temperature to 0.6 to avoid endless repetition. All inference is run on a single NVIDIA A100 GPU with 80GB VRAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
@@ 4.2 Evaluation Metrics
  • Efficiency Score: Effective communication is concise but meaningful. Inspired by the classical metric BLEU  (Papineni et al., 2002) that balances the n-gram precisions with a length penalty, we propose a new metric which balances success rate and conciseness:
    $S_{\text{conciseness}}=\frac{1}{1+\frac{\tau}{1000}},~~~S_{\text{efficiency}}=\frac{2\times\text{PSR}\times S_{\text{conciseness}}}{\text{PSR}+S_{\text{conciseness}}}$ (1)
    where $\tau$ is the average token usage per puzzle. We use the harmonic mean between the conciseness and performance scores to ensure that both need to be high to indicate good performance.
@@
4.3 Models Evaluated
We briefly summarize the types of models evaluated in our study. More details about the models can be found in Appendix Evaluated Model Details.
  • Human: We conduct experiments in which a human plays as the solver to provide an idealized upper bound. As hiring participants was prohibitively expensive and time-consuming, we role-played as agents ourselves across 100 sampled puzzles as a preliminary study, and leave further human participation to future work (see Section 7).
  • Reasoning Models: We experiment with several state-of-the-art reasoning models including o4-mini  (OpenAI, 2025), R1-OneVision  (Yang et al., 2025b), and LLaVA-CoT  (Xu et al., 2024b). We exclude o3 from evaluation, as it performs similarly to o4-mini in our exploration but incurs substantially higher computational cost.
  • General Purpose Proprietary Models: We compare with several powerful VLMs used in everyday tasks, including GPT-4o (Hurst et al., 2024), Gemini-2.0-Flash  (Anil et al., 2023), and GPT-4V  (Achiam et al., 2023).
  • General Purpose Open-Source Models: We compare with state-of-the-art open-source vision-language models pre-trained on diverse tasks, such as InternVL  (Chen et al., 2024), QwenVL  (Bai et al., 2023), LLaMA 3.2  (Touvron et al., 2024), and LLaVA 1.6  (Liu et al., 2024a).
  • Random Baseline: This baseline selects a random action uniformly from a pool of valid actions at each timestep. It does not utilize any information from the expert agent, puzzle state, conversation history, or past actions.
[…]
Table 1: Average partial success rate of multimodal agents on each puzzle with standard error estimates for uncertainty $\pm\sigma$. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column.
@@ 5.1 Overall Performance
Table 1 presents the average partial success rate (%) of various multimodal agents and a human solver across several puzzles, highlighting their relative performance. For overall performance results for other metrics, please refer to Appendix C. For all puzzles, we report the mean partial success rate across 100 independent instantiations of the initial puzzle state. Due to the source of randomness being the puzzle initialization, rather than differences in model output on the same data, we report standard error estimates for all puzzles ($\pm\sigma$) to approximate the confidence interval of performance on each puzzle.
@@
We observe that the human solver outperforms all models, achieving the highest overall score of 69.01%. Among the AI models, o4-mini demonstrates the best overall performance (53.98%), significantly surpassing the others, including GPT-4o (41.74%) and Gemini 2.0 (33.37%). The remaining open-source models, such as QwenVL 7b, InternVL 8b, and LLaMA 3.2, show significantly lower success rates, with LLaVA-CoT achieving the lowest overall score (14.97%). Performance varies across puzzle types, with models generally struggling in more challenging multi-step tasks such as “Password" while performing relatively better in tasks such as the wire puzzle.
Interestingly, the random baseline is competitive with models like LLaVA-CoT and R1-OneVision, despite having no access to the puzzle state or expert knowledge. We found that long CoT models such as LLaVA-CoT attempt to solve the puzzles independently through chain-of-thought reasoning, but they struggle to generalize to the novel communication setting introduced in our benchmark. These models are typically optimized for single-agent, instruction-following tasks, and their reasoning often fails to align with the implicit coordination required between agents. In contrast, the random baseline naturally explores a broader set of trajectories by sampling uniformly from the space of valid actions. This stochasticity allows it to occasionally stumble upon a correct solution. As a result, the random baseline can match or even outperform models that rely on self-reasoning strategies in unfamiliar collaborative contexts.
@@
Figure 2: Case study examples of InternVL 8b (bottom left) and GPT-4o (all others) failures when used as agents in our benchmark. Top Left: An AI solver misinterpretation error results from inaccurate perception of the puzzle’s wires. Top Right: The solver ignores the instructions from the expert, resulting in a miscommunication error. Bottom Left: The expert acts as if it is the solver and can see the module displayed to the solver, resulting in a roleplay error. Bottom Right: The solver performs the same action in a previously seen situation, leading to a repetition loop error.
@@
Figure 3: Distribution of error categories across conversations evaluated by our calibrated GPT-Judge for GPT-4o conversations (left), LLaMA 3.2 conversations (center), and o4-mini (right).
5.2 Qualitative Analysis on Model Failures
In this section, we present key insights and analyze common failure modes exhibited by the agents during their conversations. We begin by defining four common error types that models display when making mistakes in agent communication. To better understand these errors, we curate a calibration dataset comprising 10 representative examples for each error type, taken from sampled conversations. This dataset is used to validate and develop a o1 (Jaech et al., 2024) model judge. To calibrate the judge, we first identified common failure cases we were observing (e.g., miscommunication) that prevented the solver from finishing the puzzle or making progress. We then manually and randomly sampled 10 instances of each failure type, along with 10 randomly selected normal conversation logs. The calibration process involved continuously refining the prompt used for the GPT-Judge until it achieved high classification accuracy. After making minor adjustments to the judge’s prompt and providing access to the ground truth conversations, the model achieved 94% accuracy on the calibration dataset, demonstrating strong alignment with human annotators. Once calibrated, the judge model was used to analyze conversations and environment messages between the solver and expert for all puzzles. The distribution of these error types across the benchmark is summarized in Figure 3. We acknowledge that the GPT-Judge is not a perfect system and may exhibit some bias favoring responses that are similar to the GPT-Judge’s own patterns. We aim to mitigate this bias by ensuring that the prompt used for judgment contained clear, rule-grounded definitions of each failure type to minimize ambiguity. Moreover, the GPT-Judge is not the primary source of accuracy in our evaluation. Our primary results rely on rule-based scoring and solver progression metrics. The GPT-Judge is intended to serve as a complementary analysis tool that enables scalable evaluation of agent interactions.